# Image Corruption-Inspired Membership Inference Attacks against Large Vision-Language Models

### v2

@@ 5.1 Experimental Setup
Evaluation Metric. Following previous work Li et al. (2024b), we use Area Under the Curve (AUC) and True Positive Rate at 5% False Positive Rate (TPR at 5% FPR) as the evaluation metrics. For both metrics, higher values indicate better MIA performance. The descriptions of these metrics can be found in Appendix C.
@@
Computational Resources. We conduct all experiments on machines equipped with four NVIDIA RTX A6000 GPUs (48GB memory).
@@
Corruption Methods. We choose the following corruption methods: (i) Gaussian Blur: This is a technique that makes image soft and blurry using a Gaussian Kernel; (ii) Motion Blur: We apply a custom convolution kernel, mimicking the effect of horizontal motion-blur; and (iii) JPEG Compression: This is a method to compress the image to get the corrupted version. We apply CV2 (Bradski, 2000) to achieve Gaussian Blur and Motion Blur. We put some examples of different corruption techniques in Appendix B.
Baselines. We adopt representative and state-of-the-art MIA methods against LVLMs, including: (i) AugKL (Li et al., 2024b; Liu et al., 2021)Li et al. (2024b) extend the approach designed by Liu et al. (2021) to LVLMs. Specifically, Li et al. (2024b) quantify the difference between the original and augmented images (Such as Crop and Rotation) by computing the KL divergence between their distributions of logits; (ii) Max_Prob_Gap (Li et al., 2024b): It is the average value of the difference between the highest and second-highest token probabilities at each position. (iii) Min-K (Shi et al., 2024): Min-K is an MIA method designed for LLMs. It uses a ground-truth token and computes the lowest Kof its predicted probabilities. Li et al. (2024b) extends it to the LVLM domain; (iv) Perplexity: It is based on loss. Li et al. (2024b) analyze target perplexity to achive the attacks (Carlini et al., 2021); (v) MaxRényi-K (Li et al., 2024b): It selects the top Ktokens with the highest Rényi entropy. Then the value is the average value of these entropies; and (vi) Mod-Rényi (Li et al., 2024b): This is an extended version of MaxRényi-K which is designed for target-based scenarios.
Table 1: Image MIA AUC of various baseline methods under Li et al. (2024b)’s pipeline and our proposed approach on VL-MIA/Flickr. For all the baselines, the term "img" refers to the logits segment associated with the image embedding, while "inst" represents the instruction segment. "desp" corresponds to the generated description’s logits segment, and "inst+desp" denotes the combination of the instruction and description segments.
@@
Implementation details. All these baselines need the knowledge of the target model’s tokenizer and the output logits. The implementations of all baselines are based on Li et al. (2024b). Following the recommendation of Shi et al. (2024) and Li et al. (2024b), we set $K=20$ for the Min-K method. $K$ is set to 0, 10, and 100 for MaxRényi-$K\%$ and $\alpha$ is chosen over 0.5, 1, and 2 for both ModRényi and MaxRényi-$K\%$, as in Li et al. (2024b). We report the highest AUC for each method and provide the TPR at 5% FPR using the hyperparameter combination that achieves this highest AUC. For the Image Similarity-based attack, we fix the Kernel Size of the Gaussian blur and the Motion blur as 5. For JPEG compression, the image quality is set to 5 (lower values indicate stronger compression). For the Text Similarity-based attack, following Li et al. (2024b), we use “Describe this image concisely” (Li et al., 2024b) as the prompt and the max generation token amount is 32. For simplicity, we only use Gaussian blur for Text Similarity-based attack and the kernel size is set to 45.
Figure 6: Text similarity-based membership inference attack on VL-MIA/Flickr. Image corruption method is Gaussian Blur and the kernel size is 45.
@@ 5.2 White-Box MIA Performance
The results on LLaVA 1.5 series are shown in Table 1 and Table 2. Results on VL-MIA/Flickr 2K are in Appendix D. Note that our methods have different knowledge of the models compared to the baseline methods. We can obtain the image embeddings while they can obtain the output text’s logits. We can observe that similarity-based attacks using Gaussian Blur and Motion Blur can achieve an AUC higher than 0.8 for both LLaVA 1.5-7B and LLaVA-1.5-13B, which is much higher than all the baselines. In comparison, the highest AUC among the baseline methods is 0.743 on LLaVA-1.5-7B and 0.738 on LLaVA-1.5-13B. Similar results can be found in Table 2, our methods largely outperform all the baselines in terms of TPR at 5% FPR where the best baseline performance is 0.210 on LLaVA-1.5-7B and 0.227 on LLaVA-1.5-13B.
It shows that using information from the visual side can better facilitate image MIAs compared with using information from different logit slices. The results also suggest that the image corruption method matters. For example, the performance using JPEG compression is worse than that using Gaussian Blur and Motion Blur.
@@ 5.3 Black-box MIA performance
The results of our attack method under the black-box setting, which is based on output texts’ embeddings, are shown in Figure 6. The results demonstrate the effectiveness of our method. For example, it achieves an AUC of 0.652 and a TPR@5%FPR of 0.137 on LLaVA-1.5-7B, which outperforms many baselines that require using output logits, even though our approach only queries the target model.
5.4 Hyperparameter Analysis
In this subsection, we conduct experiments to observe the impacts of hyperparameters. For the image similarity-based attack, we vary the kernel size for Gaussian Blur and Motion Blur using values of 3, 5, 7, 15, 25, 35, and 45. The image quality for JPEG compression is selected across 1, 5, 10, 20, 45, and 70. The selected model is LLaVA 1.5 7B. The results are shown in Figure 7. We can observe that smaller kernel sizes generally contribute to better performance in terms of both AUC and TPR@5%FPR. Larger kernel sizes will bring stronger corruption. Therefore, for blur-based corruption methods, it suggests that it would be better to use a smaller corruption level for such methods. Although the performance drops a lot with very large kernel sizes, it still outperforms most baselines that rely on output logits. For the third sub-figure in Figure 7, we can see that a smaller JPEG quality (Stronger compression) generally leads to higher AUC.
For the text similarity-based attack, the kernel size for Gaussian Blur is varied across 15, 25, 35, 45 and 55. The results are shown in Figure 8. Unlike image similarity-based attacks, text similarity-based attacks generally perform better with larger kernel sizes. One possible reason is that if the kernel size is too small, small changes in the image embeddings lead to even smaller changes in the textual output. This makes the difference between member and non-member images less clear.
Method AUC TPR@5%FPR
@@ 5.5 Extra Findings
We replace the embeddings obtained by the CLIP vision encoder and the alignment module with the embeddings directly from the CLIP vision encoder without passing them through the alignment module. The results are shown in Table 3. Interestingly, we find that using the embeddings directly from the CLIP vision encoder can even have slightly better performance. However, the CLIP encoder is frozen during the training stage of LLaVA v 1.5. This suggests that the images were likely included in the CLIP model’s original training data. Our method’s high performance might also benefit from this. This suggests that we need some new benchmark datasets to better define the image MIA problem in the context of LVLMs.
@@ 6 Conclusion
In this paper, we design novel membership inference attack methods named ICIMIA against LVLMs under both white-box and black-box settings. Our approach is based on the observation that LVLMs exhibit varying sensitivity to image corruption for member and non-member images. We leverage this phenomenon by using the pair-wise similarity of the original version and its corrupted counterpart as the metric. Experimental results on existing datasets validate the effectiveness of our proposed methods.
@@
We observed an interesting phenomenon: many non-member images generated by DALL·E (Ramesh et al., 2022) exhibit greater robustness to corruption compared to their original counterparts (the generated image is prompted to be similar to its original member image), regardless of whether the original image is a member or non-member. Therefore, our method does not work for such a dataset where each non-member image is generated by DALL-E based on the original member image. One example is the img_dalle dataset Li et al. (2024b) where Blip (Li et al., 2023b) generates a caption for each member image and the caption is used by DALL-E to generate a non-member image for this image. We leave this as our future work.
@@
Our work aims to find the images that are in the training data of large vision-language models. Our proposed method, ICIMIA, can help individuals know whether their sensitive data is used to train the model. All experiments are conducted on open-source models and publicly available datasets.
AI Assistants In Writing. In this paper, we only use AI assistants for grammar checking and sentence polishing.
@@ Appendix A Details of Datasets
Many models, such as LLaVA 1.5 and Minigpt 4 (Zhu et al., 2023), use MS COCO (Lin et al., 2014) to train the models. Therefore, Li et al. (2024b) use part of the images in MS COCO as the member images. For non-member data, Li et al. (2024b) select images uploaded to Flickr [^2] https://www.flickr.com/ after those models’ release date. The datasets are licensed under the Creative Commons Attribution 4.0 International License. The dataset statistics are summarized in Table 4.
Table 4: Dataset statistics. The datasets are constructed by Li et al. (2024b)
Name Member Data Non-member Data #Memebr Data #Non-member Data
VL-MIA/Flickr MS COCO Flickr 300 300
VL-MIA/Flickr 2K MS COCO Flickr 1000 1000
Appendix B Examples of Different Corruptions
In this section, we show some examples of different types of corruption. The examples are shown in Figure 9. The original image is from the VL-MIA/Flickr (Li et al., 2024b) dataset and is sourced from: https://www.flickr.com/photos/9750464@N02/53573626031
Figure 9: Examples of different corruption methods.
Appendix C Used Metrics
@@
  • TPR at 5% FPR: True Positive Rate at 5% False Positive Rate (TPR at 5% FPR) is another widely used metric to evaluate the performance of a classification model (Li et al., 2024b; Carlini et al., 2022a). It reflects the value of the True Positive Rate when the False Positive Rate is 5%.
Appendix D Additional Results
We conduct experiments on VL-MIA/Flickr-2K. The image similarity-based MIA results are shown in Table 5 and Table 6. We can have similar observations as in Table 1 and Table 2. The experimental results on VL-MIA/Flickr-2K also validate the effectiveness of our proposed methods.
Appendix E Potential Risks
Although our proposed ICIMIA is designed to protect data integrity, malicious users can use ICIMIA to infer whether an image is used to train a certain LVLM and then get some private information. For example, if the attacker knows that one person’s medical image is used to train an LVLM for a certain disease, the attacker can get the information that this person is likely to have this disease.
Table 5: Image MIA AUC performance of our proposed approach on VL-MIA/Flickr 2K.
Method LLaVA-1.5-7B LLaVA-1.5-13B
Ours (Image_Sim, Gaussian Blur, Kernel Size 5) 0.854 0.850
Ours (Image_Sim, Motion Blur, Kernel Size 5) 0.842 0.837
Ours (Image_Sim, JPEG Compression, Quality = 5) 0.629 0.635
Table 6: Image MIA TPR at 5% FPR performance of our proposed approach on VL-MIA/Flickr 2K.
Method LLaVA-1.5-7B LLaVA-1.5-13B
Ours (Image_Sim, Gaussian Blur, Kernel Size 5) 0.349 0.337
Ours (Image_Sim, Motion Blur, Kernel Size 5) 0.372 0.352
Ours (Image_Sim, JPEG Compression, Quality = 5) 0.088 0.080

### v3

@@ 5.1 Experimental Setup
Evaluation Metric. Following previous work Li et al. (2024b), we use Area Under the Curve (AUC) and True Positive Rate at 5% False Positive Rate (TPR at 5% FPR) as the evaluation metrics. For both metrics, higher values indicate better MIA performance. The descriptions of these metrics can be found in Appendix B.
@@
Computational Resources. We conduct all experiments on machines equipped with NVIDIA RTX A6000 GPUs (48GB memory each).
@@
Corruption Methods. We choose the following corruption methods: (i) Gaussian Blur: This is a technique that makes image soft and blurry using a Gaussian Kernel; (ii) Motion Blur: We apply a custom convolution kernel, mimicking the effect of motion-blur; and (iii) JPEG Compression: This is a method to compress the image to get the corrupted version. We apply OpenCV (Bradski, 2000) to achieve Gaussian Blur and Motion Blur.
Baselines. We adopt representative and state-of-the-art MIA methods against LVLMs, including: (i) AugKL (Li et al., 2024b; Liu et al., 2021)Li et al. (2024b) extend the approach designed by Liu et al. (2021) to LVLMs. Specifically, Li et al. (2024b) quantify the difference between the original and augmented images (Such as Crop and Rotation) by computing the KL divergence between their distributions of logits; (ii) Max_Prob_Gap (Li et al., 2024b): It is the average value of the difference between the highest and second-highest token probabilities at each position; (iii) Min-K% (Shi et al., 2024): Min-K is an MIA method designed for LLMs. It uses a ground-truth token and computes the lowest K% of its predicted probabilities. Li et al. (2024b) extends it to the LVLM domain; (iv) Perplexity: It is based on loss. Li et al. (2024b) analyze target perplexity to achive the attacks (Carlini et al., 2021); (v) MaxRényi-K% (Li et al., 2024b): It selects the top K% tokens with the highest Rényi entropy. Then the value is the average value of these entropies; and (vi) Mod-Rényi (Li et al., 2024b): This is an extended version of MaxRényi-K% and is designed for target-based scenarios.
@@
Table 1: AUC of various baseline methods under Li et al. (2024b)’s pipeline and our proposed approach on VL-MIA/Flickr. For all the baselines, the term "img" refers to the logits segment associated with the image embedding, while "inst" represents the instruction segment (Li et al., 2024b). "desp" corresponds to the generated description’s logits segment, and "inst+desp" denotes the combination of the instruction and description segments (Li et al., 2024b).
@@
Implementation Details. All these baselines need the knowledge of the target model’s tokenizer and the output logits. The implementations of all baselines are based on Li et al. (2024b). Following the recommendation of Shi et al. (2024) and Li et al. (2024b), we set $K=20$ for the Min-K% method. $K$ is set to 0, 10, and 100 for MaxRényi-K% and $\alpha$ is chosen over 0.5, 1, and 2 for both ModRényi and MaxRényi-K%, as in Li et al. (2024b). We report the highest AUC for each method and provide the TPR at 5% FPR using the hyperparameter combination that achieves this highest AUC. For the Image Similarity-based attack, we fix the Kernel Size of the Gaussian blur and the Motion blur as 5. For JPEG compression, the image quality is set to 5 (lower values indicate stronger compression). For the Text Similarity-based attack, following Li et al. (2024b), we use “Describe this image concisely” (Li et al., 2024b) as the prompt and the max generation token amount is 32. For simplicity, we only use Gaussian blur for Text Similarity-based attack and the kernel size is set to 45.
Figure 5: Text similarity-based membership inference attack on VL-MIA/Flickr.
@@ 5.2 White-Box MIA Performance
The results on LLaVA 1.5 series are shown in Table 1 and Table 2. Results on VL-MIA/Flickr 2K are in Appendix C. Please note that our methods have different knowledge of the models compared with the baseline methods in Table 1 and Table 2. These baseline methods can be viewed grey-box attack methods. We can obtain the image embeddings while they can obtain the output text’s logits. We can observe that similarity-based attacks using Gaussian Blur and Motion Blur can achieve an AUC higher than 0.8 for both LLaVA 1.5-7B and LLaVA-1.5-13B, which is much higher than all the baselines. In comparison, the highest AUC among the baseline methods is 0.743 on LLaVA-1.5-7B and 0.738 on LLaVA-1.5-13B. Similar results can be found in Table 2, our methods largely outperform all the baselines in terms of TPR at 5% FPR. Our best method can have a TPR at 5% FPR higher than 0.3 on both models while the best baseline performance is 0.210 on LLaVA-1.5-7B and 0.227 on LLaVA-1.5-13B.
@@
The results show that using information from the visual side can better facilitate image MIAs compared with using information from different logit slices. The results also suggest that the choice of image corruption method plays an important role. For example, the performance using JPEG compression is worse than that using Gaussian Blur and Motion Blur.
@@ 5.3 Black-box MIA Performance
The results of our attack method under the black-box setting, which is based on output texts’ embeddings, are shown in Figure 5. The results demonstrate the effectiveness of our method. For example, it achieves an AUC of 0.652 and a TPR at 5% FPR of 0.137 on LLaVA-1.5-7B, which outperforms many baselines in Table 1 that require using output logits, even though our approach only queries the target model and utilizes the output text.
Method AUC TPR at 5% FPR
@@
5.4 Hyperparameter Analysis
In this subsection, we conduct experiments to observe the impacts of hyperparameters. For the image similarity-based attack, we vary the kernel size for Gaussian Blur and Motion Blur using values of 3, 5, 7, 15, 25, 35, and 45. The image quality for JPEG compression is selected across 1, 5, 10, 20, 45, and 70. The selected model is LLaVA 1.5 7B. The results are shown in Figure 6. We can observe that smaller kernel sizes generally contribute to better performance in terms of both AUC and TPR at 5% FPR. Larger kernel sizes will bring stronger corruption. Therefore, for blur-based corruption methods, it suggests that a smaller corruption level could have better performance. Although the performance drops a lot with very large kernel sizes, it still outperforms most baselines that rely on output logits. For the third sub-figure, we can see that a smaller JPEG quality (Stronger compression) generally leads to higher AUC.
For the text similarity-based attack, the kernel size for Gaussian Blur is varied across 15, 25, 35, 45 and 55. The results are shown in Figure 7. Unlike image similarity-based attacks, text similarity-based attacks generally perform better with larger kernel sizes. One possible reason is that if the kernel size is too small, small changes in the image embeddings lead to even smaller changes in the textual output. This makes the difference between member and non-member images less clear. Since we are using the AUC as the metric, we don’t need to explicitly set the threshold $\lambda$.
@@ 5.5 Extra Findings
We replace the embeddings obtained by the CLIP vision encoder and the alignment module with the embeddings directly from the CLIP vision encoder without passing them through the alignment module. The results are shown in Table 3. Interestingly, we find that using the embeddings directly from the CLIP vision encoder can even have slightly better performance. However, the CLIP encoder is frozen during the training stage of LLaVA-1.5. This suggests that the images might be included in the CLIP model’s original training data. Our method’s strong performance might also benefit from this. This suggests that we need some new benchmark datasets to better define the image MIA problem in the context of LVLMs.
@@ 6 Conclusion
In this paper, we design novel membership inference attack methods named ICIMIA against LVLMs under both white-box and black-box settings. Our approach is based on the observation that LVLMs exhibit varying sensitivity to image corruption for member and non-member images. We leverage this phenomenon by using the pair-wise similarity of the original version and its corrupted counterpart as the metric. Experimental results on representative datasets validate the effectiveness of our proposed methods in image membership inference attacks.
@@
We observed an interesting phenomenon: many non-member images generated by DALL·E 2 (Ramesh et al., 2022) exhibit greater robustness to corruption compared to their original counterparts (the generated image is prompted to be similar to its original member image), regardless of whether the original image is a member or non-member. Therefore, our method does not work for such a dataset where each non-member image is generated by DALL-E based on the original member image. One example is the VL-MIA/DALL-E dataset (Li et al., 2024b) where Blip (Li et al., 2023b) generates a caption for each member image and the caption is used by DALL-E to generate a non-member image for this image. We leave this as our future work.
@@
Our work aims to find the images that are in the training data of large vision-language models. Our proposed method, ICIMIA, can help individuals know whether their sensitive data is used to train the model. All experiments are conducted on open-source models and publicly available datasets. In this paper, we only use AI assistants for grammar checking and sentence polishing.
Acknowledgments
This material is based upon work supported by, or in part by the Army Research Office (ARO) under grant number W911NF-2110198, the Department of Homeland Security (DHS) under grant number 17STCIN00001-05-00, and Cisco Faculty Research Award. The findings in this paper do not necessarily reflect the view of the funding agencies.
@@ Appendix A Details of Datasets
Many models, such as LLaVA 1.5 (Liu et al., 2024a) and Minigpt 4 (Zhu et al., 2024), use MS COCO (Lin et al., 2014) to train the models. Therefore, Li et al. (2024b) use part of the images in MS COCO as the member images. For non-member data, Li et al. (2024b) select images uploaded to Flickr [^2] https://www.flickr.com/ after those target models’ release date. VL-MIA/Flickr and VL-MIA/Flickr 2K are licensed under the Creative Commons Attribution 4.0 International Public License. The dataset statistics are summarized in Table 4.
Appendix B Used Metrics
@@
  • TPR at 5% FPR: True Positive Rate at a fiex False Positive Rate is another widely used metric to evaluate the performance of MIA methods (Li et al., 2024b; Carlini et al., 2022). Following Li et al. (2024b), we use TPR at 5% FPR which reflects the value of the True Positive Rate when the False Positive Rate is 5%.
Appendix C Additional Results on More Datasets
We conduct experiments on VL-MIA/Flickr-2K. The image similarity-based MIA results are shown in Table 5. We can have similar observations as those shown in Table 1 and Table 2. The experimental results on VL-MIA/Flickr-2K also validate the effectiveness of our proposed methods.
Appendix D Potential Risks
Although our proposed ICIMIA is designed to protect data safety and privacy, malicious users can use ICIMIA to infer whether an image is used to train a certain LVLM and then get some private information. For example, if the attacker knows that one person’s medical image is used to train an LVLM for a certain disease, the attacker can get the information that this person might have this disease.